On the Apollo 14 mission to the moon, astronaut Alan Shepard hit a golf ball improvised from a tool. The free-fall acceleration on the moon is -1.63m/s^2. Suppose he hit the ball wirh a speed of 25 m/s at an angle of 30 degrees above the horizontal.

a) How long was the ball in flight
b) How far did it travel?
c) Ignoring air resistance, how much farther would it travel on the moon than the earth?
Answer #1

here,

the initial speed , u = 25 m/s

theta = 30 degree

g = - 1.63 m/s^2

a)

the time of flight for ball , t = 2 * u * sin(theta) /g

t = 2 * 25 * sin(30) /1.63 s

t = 15.3 s

b)

the distance traveled , s = u^2 * sin(2*theta) /g

s = 25^2 * sin(2*30) /1.63 m

s = 331 m

c)

the distance traveled on earth , s' = u^2 * sin(2*theta) /g

s' = 25^2 * sin(2*30) /9.81 m

s' = 55.2 m

the more distance traveled , ds = s - s' = 276 m